OpenStudy (anonymous):

@Michele_Laino A box holds 42 pieces of red construction paper and 18 pieces of blue construction paper. A piece of paper is pulled out of the box at random and then placed back in the box. How many times can you expect that a piece of red paper will be pulled out of the box in 110 tries? A. 48 B. 55 C. 60 D. 77

OpenStudy (anonymous):

I got D

OpenStudy (michele_laino):

hint: what is the probability to get a red paper?

OpenStudy (anonymous):

77?

OpenStudy (anonymous):

42/(42+18)=7/10 7/10*110=77 red papers